OpenStudy (anonymous):

Solve for x showing all of your work: 2(n + 5) = −2

OpenStudy (anonymous):

Divide both sides by 2 n+5=−2/2 Cancel 2 n+5=−1 Subtract 5 from both sides n=−1−5 Simplify −1−5 to −6 n=−6

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

yup